# On the spectral radius of the ratio of Girko matrices

### v1

@@ Abstract.
Girko matrices have independent and identically distributed entries of mean zero and unit variance. In this note, we consider the random matrix model formed by the ratio of two independent Girko matrices, its entries are dependent and heavy-tailed. Our main message is that divided by the square root of the dimension, the spectral radius of the ratio converges in distribution, when the dimension tends to infinity, to a universal heavy-tailed distribution. We provide a mathematical proof of this high-dimensional phenomenon, under a fourth moment matching with a Gaussian case known as the complex Ginibre ensemble. In this Gaussian case, the model is known as the spherical ensemble, and its spectrum is a determinantal planar Coulomb gas. Its image by the inverse stereographic projection is a rotationally invariant gas on the two-sphere. A crucial observation is the invariance in law of the model under inversion, related to its spherical symmetry, and that makes, in a sense, edge and bulk equivalent. Our approach involves Girko Hermitization, local law estimates for Wigner matrices, lower bound estimates on the smallest singular value, and convergence of kernels of determinantal point processes. The universality of the high-dimensional fluctuation of the spectral radius of the ratio of Girko matrices turns out to be remarkably more accessible mathematically than for a single Girko matrix!
@@ 1. Introduction
The asymptotic analysis in high dimension of the spectral radius of non-normal random matrices is a delicate subject, which goes back at least to Cohen and Newman (1984); Geman (1986); Bai and Yin (1986); Rider (2003). There has been recent progress in the case of Girko random matrices, see for instance Bordenave et al. (2022); Cipolloni et al. (2024a, b); Byun and Forrester (2025) and references therein. Girko matrices have independent square-integrable entries. This note addresses the case of the ratio of two independent Girko matrices, a model of non-normal random matrices with dependent entries. For this model, we prove the universality of the asymptotic behavior of the spectral radius in high dimension. This model, in the special case of Ginibre matrices, known as the spherical ensemble, becomes integrable and biunitary invariant, and the law of its spectrum is a planar heavy-tailed Coulomb gas, with a determinantal structure. The spherical ensemble is a matrix analogue and extension of the ratio of independent real normal random variables, and can be then seen as a sort of matrix Cauchy law.
@@ 1.1. Main results
Theorem 1.2 (From ratio of Girko matrices to infinite Ginibre).
Let $M=AB^{-1}$ where $A$ and $B$ are independent $n\times n$ Girko matrices, not necessarily of the same law, both satisfying (C1)(C3). Let $\mathrm{Gin}_{\infty}$ be the infinite Ginibre determinantal point process on $\mathbb{C}$, with kernel $K(z,w)=\sqrt{\gamma(z)\gamma(w)}\mathrm{e}^{z\overline{w}}$ where $\gamma(z)=\frac{1}{\pi}\mathrm{e}^{-|z|^{2}}$. Then, for all fixed $\lambda_{0}\in\mathbb{C}$, for the topology of continuous and compactly supported test functions,
@@
Note that the laws of $R_{0}=\sqrt{\min_{k\geq 1}\gamma_{k}}$ and $R_{\infty}=\displaystyle\frac{1}{R_{0}}$ have respective cdf given for all $x>0$ by
@@
where $F_{k}(x)=\mathbb{P}(\gamma_{k}<x)=\mathbb{P}(\pi_{x}<k)=\mathrm{e}^{-x}\sum_{j=0}^{k-1}\frac{x^{j}}{j!}$ with $\pi_{x}\sim\mathrm{Poisson}(x)$. In particular $R_{\infty}$ is heavy-tailed as
@@
Since ${(\xi_{k})}_{k\geq 1}$ are not i.i.d., we are beyond classical extreme value theory : $R_{0}$ is not Weibull, and $R_{\infty}$ is not Fréchet.
@@
The replacement principle stated in Theorem 1.4 also applies to the real case where $M=AB^{-1}$ with $A,B$ being real Girko matrices, and $A^{\mathrm{Gin}}$ and $B^{\mathrm{Gin}}$ being real Ginibre matrices. The proof is essentially the same with slight modifications. For the joint eigenvalue distribution of the real spherical ensemble, we refer to Forrester and Mays (2012).
@@ 1.2. Comments and open problems
The fourth moment matching in (C2) helps to reduce the complexity of the proof of Theorem 4.1, essentially the bulk universality for Girko matrices, see for instance Tao and Vu (2015). Theorem 1.3 and Theorem 4.1 are expected to remain available with the sole second moment matching.
@@ 1.2.6. Ratio of other models
We could ask about the behavior of the spectral radius of $AB^{-1}$ if $A$ and $B$ are independent real or quaternion Girko matrices, and, in another direction, real/complex/quaternion Wigner matrices.
@@ 1.3. Conventions and notation
For integers $k,l\in{\mathbb{N}}$, with $k<l$, we write $k,l]=\{k,\dots,l\}$ and $[k]=[1,k]$.
@@ 2.1. Coulomb gas and determinantal structure
$\varphi_{n}(z_{1},\ldots,z_{n})=\det\bigr[K_{n}(z_{i},z_{j})\bigr]_{1\leq i,j\leq n}.$ (2.4)
@@ 2.2. Kostlan observation and spectral radius
Moreover, we also note at this step that $n/\xi_{n,n-k}^{2}\overset{\mathrm{d}}{\to}\mathrm{Gamma}(k,1)$ as $n\to\infty$. Following Chafaï and Péché (2014); Jiang and Qi (2017), it is possible to get from (2.11) the asymptotic fluctuations of the spectral radius, namely
@@ 2.3. Equilibrium measure
$R(z)=\frac{z+\lambda_{0}}{-\overline{\lambda_{0}}z+1}.$ (3.2)
For any such rotation $R$, we can show that $|\mathrm{det}(\mathrm{d}R_{0})|=(1+|\lambda_{0}|^{2})^{2}$ because it preserves the uniform measure on the sphere $(1+|z|^{2})^{-2}\mathrm{d}^{2}z$ but using the explicit $R$ from (3.2) we can directly calculate $R^{\prime}(0)=1+|\lambda_{0}|^{2}$ so that in this case $\mathrm{d}R_{0}=1+|\lambda_{0}|^{2}$. If we prefer to use a general rotation, using that $\mathrm{d}R_{0}$ is a multiplication by a complex number of squared norm $(1+|\lambda_{0}|^{2})^{2}$ we can also write $\mathrm{d}R_{0}=(1+|\lambda_{0}|^{2})e^{\mathrm{i}\theta}$ for some real $\theta$. Finally, by Lemma 5.3, if we consider $x_{0}=0$, $\varphi_{1}=R$ and $\varphi_{2}$ as the identity map, we get that
@@
assuming the theorem holds for the origin. Since $\mathrm{Gin}_{\infty}$ is invariant under a multiplication by $e^{i\theta}$, we can forget the factor $e^{i\theta}$ so that $\mathrm{d}R_{0}(\mathrm{Gin}_{\infty})=(1+|\lambda_{0}|^{2})\mathrm{Gin}_{\infty}$. On the other hand, since $\{R(\lambda):\lambda\mbox{ is an eigenvalue of }M^{\mathrm{Gin}}\}$ has the same law as $\{\lambda:\lambda\mbox{ is an eigenvalue of }M^{\mathrm{Gin}}\}$, we get that
@@
Furthermore, the Gamma–Poisson duality gives $\mathbb{P}(\xi_{k}^{2}<x)=\mathbb{P}(\pi_{x}<k)=\mathrm{e}^{-x}\sum_{j=0}^{k-1}\frac{x^{j}}{j!}$ where $\pi_{x}\sim\mathrm{Poisson}(x)$.
@@
so that it is a consequence of the convergence of the minimum for $BA^{-1}$. Notice the immediate but interesting fact that if $A$ and $B$ have the same law, the law of the matrix $M=AB^{-1}$ is invariant in law by inversion.
@@
    for a small $\epsilon>0$, where the approximation holds in the sense of absolute mean, and $G^{z}$ is the resolvent (or Green function) of the Hermitised matrix of AzBA-zB as defined in (4.7))-(4.8) below. This step relies on the effective lower bound of the smallest singular value of Girko matrices from Jain et al. (2021) and the local law estimates for the Green function from Ajanki et al. (2014). It is worth noting that the polynomial lower bound for the smallest singular value obtained in Bordenave and Chafaï (2012) and inspired from Tao and Vu (2010) is not enough for our purposes regarding very tiny singular values, see (4.1).
@@
$\int f_{n}(x)\mathrm{d}\mu_{M}(x)=$ $\frac{1}{2\pi}\int\Delta_{z}f_{n}(z)\log|\det({\tilde{A}}{\tilde{B}}^{-1}-z)|$
@@
Recall that $H^{z}$ in (4.7) is the Hermitised matrix of $X^{z}=\widetilde{A}-z\widetilde{B}$, and $G^{z}$ is its resolvent defined in (4.8). For each fixed $z\in{\mathbb{C}}$, $X^{z}=(x_{ij})$ is still a Girko matrix with $\mathbb{E}[x_{ij}]=0$, $\mathbb{E}[(x_{ij})^{2}]=0$ and $\mathbb{E}[|x_{ij}|^{2}]=(1+|z|^{2})/n$. In particular if we set $z=0$, then $X^{z=0}$ becomes a standard Girko matrix with variance $n^{-1}$. From (Ajanki et al., 2014, Theorem 1.1) (see also Cacciapuoti et al. (2013); Bourgade et al. (2014)), fixing any small $\xi>0$, the following local laws for the resolvent $G^{z}$ with $z=0$
@@
hold uniformly for any $|w|\leq 10$ and $\eta=|\operatorname{Im}w|\geq n^{-1+\xi}$ and for any deterministic unit vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n}$, where $m_{sc}$ is the Cauchy – Stieltjes transform of the semicircle distribution $\rho_{sc}(x)=\frac{1}{2\pi}\sqrt{(4-x^{2})_{+}}$. In general for any fixed $|z|\leq C_{0}$, by a simple rescaling, the following local laws for $G^{z}$
$\Big|\big\langle\mathbf{x},\big(G^{z}(w)-m^{z}(w)\mathbf{y}\big\rangle\Big|\prec\sqrt{\frac{1}{n\eta}},\qquad\Big|\frac{1}{n}\mathrm{Tr}G^{z}(w)-m^{z}(w)\Big|\prec\frac{1}{n\eta},$ (4.17)
@@
It is easy to check that $m^{z}$ is uniformly bounded by some constant for any $|w|\leq 10$ and $|z|\leq C_{0}$. As a direct consequence of (4.17), we obtain the following estimate on the eigenvalue counting function $N(E_{1},E_{2})=\big\{\lambda_{j}\in\mathrm{Spec}(H^{z}):E_{1}\leq\lambda_{j}\leq E_{2}\big\}$:
@@
where $\rho^{z}$ is the probability distribution whose Cauchy – Stieltjes transform is given by (4.18). The proof of (4.19) is quite standard (see similar arguments used to prove (Benaych-Georges and Knowles, 2017, Theorem 2.8)), so we omit it. We remark that though the above bound holds with very high probability for each fixed $z$, they can be extended to very high probability estimates simultaneously for all $|z|\leq C_{0}$ by a grid argument and union bound, since $G^{z}$ is Lipschitz continuous in $z$ and $\|G^{z}(\mathrm{i}\eta)\|\leq\eta^{-1}$. Without specific mention, all the estimates below hold simultaneously for all $|z|\leq C_{0}$.
Now we are ready to prove Lemma 4.2. By a simple spectral decomposition of $H^{z}$, we have
@@
Define the following the Ornstein – Uhlenbeck matrix flow
$\mathrm{d}{\tilde{A}}_{t}=-\frac{1}{2}{\tilde{A}}_{t}\mathrm{d}t+\frac{\mathrm{d}\mathcal{M}^{(1)}_{t}}{\sqrt{n}},\qquad\mathrm{d}{\tilde{B}}_{t}=-\frac{1}{2}{\tilde{B}}_{t}\mathrm{d}t+\frac{\mathrm{d}\mathcal{M}^{(2)}_{t}}{\sqrt{n}},$ (4.29)
@@
with initial ensembles ${\tilde{A}}_{0}={\tilde{A}}=A/\sqrt{n}$ and ${\tilde{B}}_{0}={\tilde{B}}=B/\sqrt{n}$, where $A,B$ are independent Girko matrices satisfying (C1)(C3), and $\mathcal{M}^{(1)}_{t}$ and $\mathcal{M}^{(2)}_{t}$ are two independent $n\times n$ matrices whose entries are i.i.d. standard complex Brownian motions. Then for any $z\in{\mathbb{C}}$ we define $X^{z}_{t}={\tilde{A}}_{t}-z{\tilde{B}}_{t}$ as in (4.6). In distribution, it satisfies

### v2

@@ Abstract.
Girko matrices have independent and identically distributed entries of mean zero and unit variance. In this note, we consider the random matrix model formed by the ratio of two independent Girko matrices. Such a random matrix is heavy tailed with dependent entries. Our main message is that divided by the square root of the dimension, the spectral radius of the ratio converges in distribution, when the dimension tends to infinity, to a universal heavy-tailed distribution. We provide a mathematical proof of this high-dimensional phenomenon, under a fourth moment matching with a Gaussian case known as the complex Ginibre ensemble. In this Gaussian case, the model is known as the spherical ensemble, and its spectrum is a determinantal planar Coulomb gas. Its image by the inverse stereographic projection is a rotationally invariant gas on the two-sphere. A crucial observation is the invariance in law of the model under inversion, related to its spherical symmetry, and that makes, in a sense, edge and bulk equivalent. Our approach involves Girko Hermitization, local law estimates for Wigner matrices, lower bound estimates on the smallest singular value, and convergence of kernels of determinantal point processes. The universality of the high-dimensional fluctuation of the spectral radius of the ratio of Girko matrices turns out to be remarkably more accessible mathematically than for a single Girko matrix!
@@ 1. Introduction
The asymptotic analysis in high dimension of the spectral radius of non-normal random matrices is a delicate subject, which goes back at least to Cohen and Newman (1984); Geman (1986); Bai and Yin (1986); Rider (2003). There has been recent progress in the case of Girko random matrices, see for instance Bordenave et al. (2022); Cipolloni et al. (2024a, b); Byun and Forrester (2025) and references therein. Girko matrices have independent square-integrable entries. This note addresses the case of the ratio of two independent Girko matrices, a model of heavy-tailed non-normal random matrices. For this model, we prove the universality of the asymptotic behavior of the spectral radius in high dimension. This model, in the special case of Ginibre matrices, known as the spherical ensemble, becomes integrable and biunitary invariant, and the law of its spectrum is a planar heavy-tailed Coulomb gas with a determinantal structure. The spherical ensemble is a matrix analogue and extension of the ratio of independent real Gaussian random variables, and can be then seen as a sort of matrix Cauchy law.
@@ 1.1. Main results
This note is concerned with the space of point configurations endowed with the weakest topology such that the maps $\mu_{\mathcal{X}}\mapsto\int f\mathrm{d}\mu_{\mathcal{X}}$ are continuous for every continuous compactly supported function $f$. As a consequence, the convergence in distribution for a sequence of point processes is always taken with respect to this topology.
Theorem 1.2 (From ratio of Girko to infinite Ginibre).
Let $M=AB^{-1}$ where $A$ and $B$ are independent $n\times n$ Girko matrices, not necessarily of the same law, both satisfying (C1)(C3). Let $\mathrm{Gin}_{\infty}$ be the infinite Ginibre determinantal point process on $\mathbb{C}$, with kernel $K(z,w)=\sqrt{\gamma(z)\gamma(w)}\mathrm{e}^{z\overline{w}}$ where $\gamma(z)=\frac{1}{\pi}\mathrm{e}^{-|z|^{2}}$. Then, for all fixed $\lambda_{0}\in\mathbb{C}$,
@@
Remark 1.4 (Tail).
The laws of $R_{0}=\sqrt{\min_{k\geq 1}\gamma_{k}}$ and $R_{\infty}=\displaystyle\frac{1}{R_{0}}$ have respective cdf given for all $x>0$ by
@@
where $1-F_{k}(x)=\mathbb{P}(\gamma_{k}\geq x)=\mathbb{P}(\pi_{x}<k)=\mathrm{e}^{-x}\sum_{j=0}^{k-1}\frac{x^{j}}{j!}$ with $\pi_{x}\sim\mathrm{Poisson}(x)$. In particular $R_{\infty}$ is heavy-tailed as
@@
We are beyond classical extreme value theory: ${(\xi_{k})}_{k\geq 1}$ are not i.i.d., $R_{0}$ is not Weibull, and $R_{\infty}$ is not Fréchet.
Remark 1.5 (Point process of farthest particles).
Actually, the proof of Theorem 1.3 gives much more: we not only obtain the limit of the farthest particle, but also of the second farthest particle, the third and so on. We may formulate this as saying that for every continuous function $f:\mathbb{C}\to\mathbb{R}$ whose support does not contain the origin,
$\sum_{\lambda\in\mathrm{spec}(M)}f(\lambda/\sqrt{n})\xrightarrow[n\to\infty]{\mathrm{d}}\sum_{w\in\mathrm{Gin}_{\infty}}f(1/w).$
Since we do not ask $f$ to be compactly supported, we are able to obtain the information of the farthest particles.
@@
Though lacking analogous versions of Theorems 1.2-1.3 for the real spherical ensemble, the replacement principle stated in Theorem 1.6 applies to the real case where $M=AB^{-1}$ with $A,B$ being real Girko matrices, and $A^{\mathrm{Gin}}$ and $B^{\mathrm{Gin}}$ being real Ginibre matrices. The proof is essentially the same, with slight modifications. Indeed, the eigenvalues of the real spherical ensemble can be studied using Pfaffians instead of determinants as explained in Forrester and Mays (2012).
@@ 1.2. Comments and open problems
The fourth moment matching in (C2) helps to reduce the complexity of the proof of Theorem 4.1, essentially for the bulk universality for Girko matrices, see for instance Tao and Vu (2015). Since bulk universality for Girko matrices with a second moment matching was proved recently in Maltsev and Osman (2024), we expect that Theorem 1.3 also holds under the assumption of second moment matching only.
@@ 1.2.6. Ratio of other models
We could ask about the behavior of the spectral radius of $AB^{-1}$ if $A$ and $B$ are independent real or quaternion Girko matrices, and, in another direction, real/complex/quaternion Wigner matrices. As mentioned in Remark 1.7, the replacement principle is expected to be true with slight modifications, though the statements for the ratios of Ginibre/Gaussian ensembles are missing.
@@ 1.3. Conventions and notation
For integers $k,l\in{\mathbb{N}}$, with $k<l$, we write $k,l]]=\{k,\dots,l\}$ and $[[k]]=[[1,k]]$.
@@ 2.1. Coulomb gas and determinantal structure
$\varphi_{n}(z_{1},\ldots,z_{n})=\frac{1}{n!}\det\bigr[K_{n}(z_{i},z_{j})\bigr]_{1\leq i,j\leq n}.$ (2.4)
@@ 2.2. Kostlan observation and spectral radius
Moreover, we also note at this step that $n/\xi_{n,n-k+1}^{2}\overset{\mathrm{d}}{\to}\mathrm{Gamma}(k,1)$ as $n\to\infty$. Following Chafaï and Péché (2014); Jiang and Qi (2017), it is possible to get from (2.11) the asymptotic fluctuations of the spectral radius, namely
@@ 2.3. Equilibrium measure
$R(z)=\frac{z+\lambda_{0}}{-\overline{\lambda_{0}}z+1}$ (3.2)
and let us denote $\mathrm{d}R_{0}$ the differential of $R$ at the point 0. If we use the explicit matrix from (3.2) we can directly calculate the complex derivative $R^{\prime}(0)=1+|\lambda_{0}|^{2}$ so that $\mathrm{d}R_{0}=(1+|\lambda_{0}|^{2}){\mathrm{Id}}_{2}$. If we prefer to use a general rotation $R$, we can show that $|\det(\mathrm{d}R_{0})|=(1+|\lambda_{0}|^{2})^{2}$ because it preserves the uniform measure on the sphere $(1+|z|^{2})^{-2}d^{2}z$, and using that $\mathrm{d}R_{0}$ is a multiplication by a complex number of squared norm $(1+|\lambda_{0}|^{2})^{2}$ we have $\mathrm{d}R_{0}=(1+|\lambda_{0}|^{2})\mathrm{e}^{\mathrm{i}\theta}$ for some real $\theta$. Finally, by Lemma 5.3, if we consider $x_{0}=0$, $\varphi_{1}=R$ and $\varphi_{2}$ as the identity map, we get that
@@
assuming the theorem holds for the origin. Since $\mathrm{Gin}_{\infty}$ is invariant under a multiplication by $\mathrm{e}^{i\theta}$, we can forget the factor $\mathrm{e}^{i\theta}$ so that $\mathrm{d}R_{0}(\mathrm{Gin}_{\infty})=(1+|\lambda_{0}|^{2})\mathrm{Gin}_{\infty}$. On the other hand, since $\{R(\lambda):\lambda\mbox{ is an eigenvalue of }M^{\mathrm{Gin}}\}$ has the same law as $\{\lambda:\lambda\mbox{ is an eigenvalue of }M^{\mathrm{Gin}}\}$, we get that
@@
Furthermore, the Gamma–Poisson duality gives $\mathbb{P}(\xi_{k}^{2}>x)=\mathbb{P}(\pi_{x}<k)=\mathrm{e}^{-x}\sum_{j=0}^{k-1}\frac{x^{j}}{j!}$ where $\pi_{x}\sim\mathrm{Poisson}(x)$.
@@
so that it is a consequence of the convergence of the minimum for $BA^{-1}$.
@@
    for a small $\epsilon>0$, where the approximation holds in the sense of absolute mean, and $G^{z}$ is the resolvent (or Green function) of the Hermitised matrix of AzBA-zB as defined in (4.7)-(4.8) below. This step relies on the effective lower bound of the smallest singular value of Girko matrices from Jain et al. (2021) and the local law estimates for the Green function from Ajanki et al. (2014). It is worth noting that the polynomial lower bound for the smallest singular value obtained in Bordenave and Chafaï (2012) and inspired from Tao and Vu (2010) is not enough for our purposes regarding very tiny singular values, see (4.1).
@@
$\int f_{n}(x)\mathrm{d}\mu_{M}(x)=$ $\frac{1}{2\pi}\int\Delta_{z}f_{n}(z)\log|\det({\tilde{A}}{\tilde{B}}^{-1}-z)|\mathrm{d}^{2}z$
@@
Recall that $H^{z}$ in (4.7) is the Hermitised matrix of $X^{z}=\widetilde{A}-z\widetilde{B}$, and $G^{z}$ is its resolvent defined in (4.8). For each fixed $z\in{\mathbb{C}}$, $X^{z}=(x_{ij})$ is still a Girko matrix with $\mathbb{E}[x_{ij}]=0$, $\mathbb{E}[(x_{ij})^{2}]=0$ and $\mathbb{E}[|x_{ij}|^{2}]=(1+|z|^{2})/n$. In particular if we set $z=0$, then $X^{z=0}$ becomes a standard Girko matrix with variance $n^{-1}$. From (Ajanki et al., 2014, Theorem 1.1) (see also Knowles and Yin (2013); Cacciapuoti et al. (2013); Bourgade et al. (2014)), fixing any small $\xi>0$, the following local laws for the resolvent $G^{z}$ with $z=0$
@@
hold uniformly for any $|w|\leq 10$ and $\eta=|\operatorname{Im}w|\geq n^{-1+\xi}$ and for any deterministic unit vectors $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n}$, where $m_{sc}$ is the Cauchy – Stieltjes transform of the semicircle distribution $\rho_{sc}(x)=\frac{1}{2\pi}\sqrt{(4-x^{2})_{+}}$. Note that for each fixed $z\in{\mathbb{C}}$, the independence of $A$ and $B$ implies that $H^{z}$ has the same distribution as $\sqrt{1+|z|^{2}}\tilde{H}$, where $\tilde{H}$ is the Hermitised matrix obtained from a standard Girko matrix with variance $1/n$ (note that the density of $\tilde{H}$ could depend on $z$). Thus, the resolvent of $\tilde{H}$ satisfies the same local law as in (4.16). A simple rescaling then yields, for each fixed $z\in{\mathbb{C}}$, the following local laws for $G^{z}$,
$\Big|\big\langle\mathbf{x},\big(G^{z}(w)-m^{z}(w)\big)\mathbf{y}\big\rangle\Big|\prec\sqrt{\frac{1}{n\eta}},\qquad\Big|\frac{1}{n}\mathrm{Tr}G^{z}(w)-m^{z}(w)\Big|\prec\frac{1}{n\eta},$ (4.17)
@@
It is easy to check that $m^{z}$ is uniformly bounded by some constant for any $|w|\leq 10$ and $|z|\leq C_{0}$. We remark that though the above bound holds with very high probability for each fixed $z$, they can be extended to very high probability estimates simultaneously for all $|z|\leq C_{0}$ by a grid argument and union bound, since $G^{z}$ is Hölder continuous in $z$ and $\|G^{z}(\mathrm{i}\eta)\|\leq\eta^{-1}$. Without specific mention, all the estimates below hold simultaneously for all $|z|\leq C_{0}$. Moreover, as a direct consequence of (4.17), we obtain the following estimate on the eigenvalue counting function $N(E_{1},E_{2})=\big\{\lambda_{j}\in\mathrm{Spec}(H^{z}):E_{1}\leq\lambda_{j}\leq E_{2}\big\}$:
@@
where $\rho^{z}$ is the probability distribution whose Cauchy – Stieltjes transform is given by (4.18). The proof of (4.19) is quite standard (see similar arguments used to prove (Benaych-Georges and Knowles, 2017, Theorem 2.8)), so we omit it.
Now we are ready to prove Lemma 4.2. The argument is similar to the proof of (Cipolloni et al., 2021, Lemma 4). By a simple spectral decomposition of $H^{z}$, we have
@@
Define the following the Ornstein – Uhlenbeck matrix flows
$\mathrm{d}{\tilde{A}}_{t}=-\frac{1}{2}{\tilde{A}}_{t}\mathrm{d}t+\frac{\mathrm{d}\mathcal{B}^{(1)}_{t}}{\sqrt{n}},\qquad\mathrm{d}{\tilde{B}}_{t}=-\frac{1}{2}{\tilde{B}}_{t}\mathrm{d}t+\frac{\mathrm{d}\mathcal{B}^{(2)}_{t}}{\sqrt{n}},$ (4.29)
@@
with initial ensembles ${\tilde{A}}_{0}={\tilde{A}}=A/\sqrt{n}$ and ${\tilde{B}}_{0}={\tilde{B}}=B/\sqrt{n}$, where $A,B$ are independent Girko matrices satisfying (C1)(C3), and $\mathcal{B}^{(1)}_{t}$ and $\mathcal{B}^{(2)}_{t}$ are two independent $n\times n$ matrices whose entries are i.i.d. standard complex Brownian motions. Then for any $z\in{\mathbb{C}}$ we set $X^{z}_{t}={\tilde{A}}_{t}-z{\tilde{B}}_{t}$ as in (4.6), use $X^{z}_{t}$ to construct the Hermitised matrix $H^{z}_{t}$ as in (4.7) and define the corresponding Green function as in (4.8), i.e.,
$H^{z}_{t}=\begin{pmatrix}0&X_{t}^{z}\\ (X_{t}^{z})^{*}&0\end{pmatrix},\quad\qquad G^{z}_{t}(w)=(H^{z}_{t}-w)^{-1},\qquad w\in{\mathbb{C}}\setminus{\mathbb{R}},\quad z\in{\mathbb{C}},\quad t\geq 0.$ (4.30)
We additionally introduce the following notations:
$H_{t}^{(1)}=\begin{pmatrix}0&{\tilde{A}}_{t}\\ ({\tilde{A}}_{t})^{*}&0\end{pmatrix},\qquad H_{t}^{(2)}=\begin{pmatrix}0&{\tilde{B}}_{t}\\ ({\tilde{B}}_{t})^{*}&0\end{pmatrix},\qquad\mathrm{d}\mathcal{M}^{(l)}_{t}=\begin{pmatrix}0&\mathrm{d}\mathcal{B}^{(l)}_{t}\\ \big(\mathrm{d}\mathcal{B}^{(l)}_{t}\big)^{*}&0\end{pmatrix},\qquad l=1,2.$ (4.31)
For notational simplicity, we may often drop the dependence $t$ and write $H^{(l)}_{t}=(h^{(l)}_{ij})_{i,j\in[[2n]]}$ for $l=1,2$. Note that the non-zero entries $(h^{(l)}_{aB})$ for $a\in[[n]]$ and $B\in[[n+1,2n]]$ are independent random variables that depend on time.
Recall the definition of $I^{\geq\eta_{0}}$ in (4.13). Note that $\mathrm{Tr}G_{t}^{z}(\mathrm{i}\eta)$ is purely imaginary due to the block structure of $H_{t}^{z}$ in (4.30), so we have $\mathrm{Tr}G_{t}^{z}(\mathrm{i}\eta)=\mathrm{i}\operatorname{Im}\mathrm{Tr}G_{t}^{z}(\mathrm{i}\eta)$. Then we define the following short-hand
with $T=n^{100}$ and $\eta_{0}=n^{-1-\epsilon}$ for $0<\epsilon<c_{0}/100$. We then use the matrix flows (4.29) and Itô’s formula to derive the dynamics of $G^{z}_{t}$, and hence $I^{\geq\eta_{0}}_{t}$, as a function of $\big\{h^{(l)}_{aB}\big\}$ in (4.31) for $a\in[[n]]$ and $B\in[[n+1,2n]]$. More precisely,
$\mathrm{d}I^{\geq\eta_{0}}_{t}=$ $\Theta\mathrm{d}t+\mathrm{d}M_{t},$ (4.33)
with the drift function $\Theta$ given by
$\Theta=\sum_{a=1}^{n}\sum_{B=n+1}^{2n}\sum_{l=1}^{2}\left(-\frac{1}{2}h^{(l)}_{aB}\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}}-\frac{1}{2}\overline{h^{(l)}_{aB}}\frac{\partial I^{\geq\eta_{0}}_{t}}{\partial\overline{h^{(l)}_{aB}}}+\frac{1}{n}\frac{\partial^{2}I^{\geq\eta_{0}}_{t}}{\partial h^{(l)}_{aB}\partial\overline{h^{(l)}_{aB}}}\right)$=: $\Theta_{1}+\Theta_{2}+\Theta_{3},$ (4.34)
and the diffusion term $\mathrm{d}M_{t}$ given by